OpenStudy (anonymous):

Quick Question! Medal and fan Calculate the discriminant to determine the number of real roots. y = x2 – 6x + 9. How many real roots does the equation have? A.one real root B.two real roots C.no real roots D.no solution to the equation

OpenStudy (anonymous):

@dmndlife24 @Ilovecake @Thesmarterone

OpenStudy (er.mohd.amir):

discriminant=b*b-4*a*c here a=1,b=-6,c=9

OpenStudy (er.mohd.amir):

if d is positive then then two real roots and zero then two roots are same that is one root which is equal to other.

OpenStudy (whpalmer4):

If the discriminant (\(\Delta\)) is 0, the equation has a single real root (with multiplicity 2). If the discriminant is > 0, the equation has two real roots. If the discriminant is < 0, the equation has two complex roots. Remember, the discriminant is the quantity under the square root sign in the equation for the solutions of a quadratic: \[x =\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

hmmm...okay, i think i understand, not very sure though